DBOY:

When simplified, which of the following expressions has a coefficient of 5? -4 x - 9 x 4 x - (-9 x) -4 x + 9 x 4 x - 9 x

Salmon:

You need to find out what plus what is 5 -4x-9x = -13x 4x+9x = 13x -4x+9x=? 4x-9x= --5x